# Nonlocal pseudo-symmetries and Bäcklund transformations as $\mathcal{C}$-morphisms

### v2

@@ 2.4 Differentiable coverings and Bäcklund transformations as nonlocal $\mathcal{C}$-morphisms
$\left\{\begin{array}[]{l}\rho_{x}=-\rho^{2}-z-\lambda,\vspace{5pt}\\ \rho_{t}=-2(2\lambda-z)\rho^{2}-2z_{x}\rho+z_{xx}+2z^{2}-2\lambda z-4\lambda^{2},\vspace{5pt}\\ z_{t}=-6zz_{x}-z_{xxx}.\end{array}\right.$ (26)
In particular, $\mu:=\rho\,dx+\left(z_{x}+\left(4\lambda-2z\right)\rho\right)dt$ is a nonlocal conservation law (depending on the nonlocal function $\rho$ defined in $\tau$). Thus, the $2$-dimensional covering can be seen as an extension of $\tau$ by means of the equations for the potential $\ln v^{1}$ of $\Lambda$. We also notice that, in terms of the Lax “potential” $\phi$, the nonlocal variable of $\tau$ reads $\rho=\phi_{x}/\phi$.
@@
Let $\mathcal{E}\subset J^{k}(\pi)$ and $\mathcal{E}^{\prime}\subset J^{k^{\prime}}(\pi^{\prime})$ be two formally integrable equations, with $\pi:E\rightarrow M$ and $\pi^{\prime}:E^{\prime}\rightarrow M^{\prime}$ such that $\dim M=\dim M^{\prime}$. A Bäcklund transformation (BT) from $\mathcal{E}$ to $\mathcal{E}^{\prime}$ is a regular $\mathcal{C}$-morphism $\mathcal{B}:J^{\infty}(\pi)\rightarrow J^{\infty}(\pi^{\prime})$ from $\mathcal{Y}$ to $\mathcal{E}^{\prime}$, for some differentiable covering $\tau:\mathcal{Y}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$, with $\mathcal{Y}\subset J^{l}(\pi)$. In particular $\mathcal{B}$ is an auto-Bäcklund transformation (ABT) whenever $\mathcal{E}^{\prime}$ is a copy of $\mathcal{E}$.
According to this definition, a BT can be seen as a pair of $\mathcal{C}$-morphisms
$\begin{array}[]{ccc}&\mathcal{Y}^{(\infty)}\\ \qquad\tau\swarrow\vspace{3pt}&&\searrow\mathcal{B}\vspace{3pt}\\ \mathcal{E}^{(\infty)}&&\qquad\mathcal{E}^{\prime}{}^{(\infty)}\end{array}$
which is analogous to the point of view followed by Krasil’shchik and Vinogradov in [23]. However, in this article, we will not limit ourselves to considering Bäcklund transformations that keep the independent variables fixed.
@@
Continuing Example 16 we can provide an example of auto-Bäcklund transformation (ABT) for KdV. Indeed, by considering the trivial bundle $\pi:\mathbb{R}^{4}\rightarrow\mathbb{R}^{2}$, $(x,t,z,\rho)\mapsto(x,t)$ and the $1$-dimensional covering $\tau:\mathcal{Y}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$, it can be checked that the $\mathcal{C}$-morphism $\mathcal{B}:J^{\infty}(\pi)\rightarrow J^{\infty}(\pi^{\prime})$ defined by
@@
through the prolongation formulas (5), defines a $\mathcal{C}$-morphism from $\mathcal{Y}$ to =$\mathcal{E}^{\prime}=\{z^{\prime}_{t}+6z^{\prime}z^{\prime}_{x}+z^{\prime}_{xxx}=0\}$, which is a copy of $\mathcal{E}$. Hence (27) provides an ABT of KdV. In practice, for any known solution $f_{0}$ of the KdV, such a transformation associates a new solution
@@
where $g_{0}$ is the solution of the Riccati system corresponding to $f_{0}$ for some $\lambda_{0}$, i.e.,
$\left\{\begin{array}[]{l}g_{0x}=-g_{0}^{2}-f_{0}-\lambda_{0},\vspace{5pt}\\ g_{0t}=-2(2\lambda_{0}-f_{0})g_{0}^{2}-2f_{0x}g_{0}+f_{0xx}+2f_{0}^{2}-2\lambda_{0}f_{0}-4\lambda_{0}^{2},\vspace{5pt}\end{array}\right.$
@@
$f_{i}=-f_{i-1}-2g_{i-1}^{2}-2\lambda_{i-1}$
where each $g_{i-1}$ is the solution of the Riccati system corresponding to $f_{i-1}$ and $\lambda_{i-1}$.
@@
Since $\rho=\phi_{x}/\phi$, one can readily check that in terms of the Lax potential above ABT can be written as
$z^{\prime}=z+2\left(\ln\phi\right)_{xx}.$
Thus, for any known solution $f_{0}$ of the KdV, the transformation associates the new solution $f_{1}=f_{0}+2\left(\ln g_{0}\right)_{xx}$ where $g_{0}$ is the corresponding solution of Lax system
$\left\{\begin{array}[]{l}g_{0xx}=-f_{0}g_{0}-\lambda_{0}g_{0},\vspace{5pt}\\ g_{0t}=2(2\lambda_{0}-f_{0})g_{0x}+f_{0x}g_{0}.\vspace{5pt}\end{array}\right.$
@@
If in the previous example we augment the space by passing to the bundle $\hat{\pi}:\mathbb{R}^{5}\rightarrow\mathbb{R}^{2},(x,t,z,\rho,\hat{\rho})\mapsto(x,t)$, and instead of (26) consider the system $\hat{\mathcal{Y}}$
$\left\{\begin{array}[]{l}\hat{\rho}_{x}=-\hat{\rho}^{2}-z-\hat{\lambda},\vspace{5pt}\\ \hat{\rho}_{t}=-2(2\hat{\lambda}-z)\hat{\rho}^{2}-2z_{x}\hat{\rho}+z_{xx}+2z^{2}-2\hat{\lambda}z-4\hat{\lambda}^{2},\vspace{5pt}\\ \rho_{x}=-\rho^{2}-z-\lambda,\vspace{5pt}\\ \rho_{t}=-2(2\lambda-z)\rho^{2}-2z_{x}\rho+z_{xx}+2z^{2}-2\lambda z-4\lambda^{2},\vspace{5pt}\\ z_{t}=-6zz_{x}-z_{xxx}.\end{array}\right.$ (28)
obtained by doubling the Riccati system, i.e., by augmenting another copy of the Riccati system with spectral parameter $\hat{\lambda}\neq\lambda$, then the natural projection $\hat{\tau}:\hat{\mathcal{Y}}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$ is another covering of KdV and the $\mathcal{C}$-morphism $J^{\infty}(\hat{\pi})\rightarrow J^{\infty}(\pi^{\prime})$ defined by
@@
is a $\mathcal{C}$-morphism from $\hat{\mathcal{Y}}$ to the system $\mathcal{E}^{\prime}$ defined by
$\left\{\begin{array}[]{l}\rho^{\prime}_{x}=-\rho^{\prime}{}^{2}-z^{\prime}-\lambda,\vspace{5pt}\\ \rho^{\prime}_{t}=-2(2\lambda-z^{\prime})\rho^{\prime}{}^{2}-2z^{\prime}_{x}\rho^{\prime}+z^{\prime}_{xx}+2z^{\prime}{}^{2}-2\lambda z^{\prime}-4\lambda^{2},\vspace{5pt}\\ z^{\prime}_{t}=-6z^{\prime}z^{\prime}_{x}-z^{\prime}_{xxx}.\end{array}\right.$
In this case, to obtain a new solution $z^{\prime}$ of KdV out of a given one $z=f_{0}$, the first step is to compute the corresponding solution $\rho=g_{0}$ of the Riccati system, with arbitrary $\lambda$. Then, since $g_{0}$ depends on $\lambda$, one gets another instance $\hat{g}_{0}$ of that function, by substituting $\lambda$ with $\hat{\lambda}\neq\lambda$. This way (29) provides the new pair of functions
$z^{\prime}=-f_{0}-2\hat{g}_{0}^{2}-2\hat{\lambda},\qquad\rho^{\prime}=-\frac{\hat{\lambda}-\lambda}{\hat{g}_{0}-g}-\hat{g}_{0}$
@@
$\left\{\begin{array}[]{l}\rho_{x}=-\rho^{2}-z-\lambda,\vspace{5pt}\\ \rho_{t}=-2(2\lambda-z)\rho^{2}-2z_{x}\rho+z_{xx}+2z^{2}-2\lambda z-4\lambda^{2},\vspace{5pt}\\ z_{t}=-6zz_{x}-z_{xxx}.\end{array}\right.$
The peculiarity of this transformation, in comparison with that of previous example, is that it provides the new solution $z^{\prime}$ together with the corresponding solution $\rho^{\prime}$ of the Riccati system, with arbitrary $\lambda$. Thus, further applications of the transformations only require derivations and algebraic operations.
@@
which is known in literature as the Darboux transformation of KdV (see for instance [19, 30]). Thus, the notion of BT described by Definition 17 allows one to regard also Darboux transformations as examples of BT.
@@
Consider the covering $\tau:\mathcal{Y}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$ of the short-pulse equation =$\mathcal{E}=\{z_{xt}=\frac{1}{2}z^{2}z_{xx}+zz_{x}^{2}+z\}\subset J^{3}(\pi)$, defined by the system $\mathcal{Y}$
$\left\{\begin{array}[]{l}\rho_{x}=-{\displaystyle\frac{z_{x}}{4\eta}}\rho^{2}+{\displaystyle\frac{1}{2\eta}}\,\rho+{\displaystyle\frac{z_{x}}{4\eta}},\vspace{10pt}\\ \rho_{t}=-{\displaystyle\frac{z\left(zz_{x}+4\,\eta\right)}{8\eta}}\rho^{2}+{\displaystyle\frac{\left(8\,\eta^{2}+z^{2}\right)}{4\eta}}\rho-{\displaystyle\frac{z\left(-zz_{x}+4\,\eta\right)}{8\eta}},\vspace{10pt}\\ z_{xt}={\displaystyle\frac{1}{2}}z^{2}z_{xx}+zz_{x}^{2}+z,\end{array}\right.$ (30)
where $\pi:\mathbb{R}^{4}\rightarrow\mathbb{R}^{2}$, $(x,t,z,\rho)\mapsto(x,t)$. One can check that the transformation
$x^{\prime}=x+8\frac{\eta}{\rho^{2}+1},\qquad t^{\prime}=t,\qquad z^{\prime}=-z+8\eta\frac{\rho}{\rho^{2}+1}.$ (31)
determines an auto-Bäcklund transformation, i.e., a BT from =$\mathcal{E}=\{z_{xt}=\frac{1}{2}z^{2}z_{xx}+zz_{x}^{2}+z\}$ to =$\mathcal{E}^{\prime}=\{z^{\prime}_{xt}=\frac{1}{2}z^{\prime}{}^{2}z^{\prime}_{xx}+z^{\prime}z^{\prime}{}_{x}^{2}+z^{\prime}\}$.

### v5

@@ 2.4 Differentiable coverings and Bäcklund transformations as nonlocal $\mathcal{C}$-morphisms
$\mathcal{Y}:\quad\left\{\begin{array}[]{l}\rho_{x}=-\rho^{2}-z-\lambda,\vskip 5.0pt\\ \rho_{t}=-2(2\lambda-z)\rho^{2}-2z_{x}\rho+z_{xx}+2z^{2}-2\lambda z-4\lambda^{2},\vskip 5.0pt\\ z_{t}=-6zz_{x}-z_{xxx},\end{array}\right$. (27)
by means of the two differential equations $\left(\ln v^{1}\right)_{x}=\rho$ and $\left(\ln v^{1}\right)_{t}=z_{x}+\left(4\lambda-2z\right)\rho$, for the potential $\ln v^{1}$ of the conservation law $\mu:=\rho\,dx+\left(z_{x}+\left(4\lambda-2z\right)\rho\right)dt$, defined in $\tau$.
Example 17.
The Tzitzeica equation $\mathcal{E}=\{z_{xt}=-{{\rm e}^{-2\,z}}+{{\rm e}^{z}}\}$, with $z=z(x,t)$, admits the ZCR $\alpha=X\,dx+T\,dt$ given by [9]
$X=\left(\begin{array}[]{ccc}-z_{x}&0&\lambda\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\lambda&z_{x}&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 0&\lambda&0\end{array}\right),\qquad T=\left({\displaystyle\begin{array}[]{ccc}0&\frac{1}{\lambda}{{\rm e}^{-2\,z}}&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 0&0&\frac{1}{\lambda}{{\rm e}^{z}}\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\frac{1}{\lambda}{{\rm e}^{z}}&0&0\end{array}}\right),$
with $\lambda$ a real parameter. Thus, by introducing the fiber bundle $\mathbb{R}^{2}\times\mathbb{R}^{4}\rightarrow\mathbb{R}^{2}$, $(x,t,z,v^{1},v^{2},v^{3})\mapsto(x,t)$, the infinite prolongation of the linear system
$\mathcal{V}:\quad\left\{\begin{array}[]{lll}v_{x}^{1}=-z_{x}v^{1}+\lambda v^{3},&\vskip 5.0pt&{\displaystyle v_{t}^{1}=\frac{1}{\lambda}e^{-2z}v^{2}},\\ v_{x}^{2}=\lambda v^{1}+z_{x}v^{2},&\vskip 5.0pt&{\displaystyle v_{t}^{2}=\frac{1}{\lambda}e^{z}v^{3}},\\ v_{x}^{3}=\lambda v^{2},&&{\displaystyle v_{t}^{3}=\frac{1}{\lambda}e^{z}v^{1}},\end{array}\right$. (28)
defines a $3$-dimensional differentiable covering $\mathcal{V}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$ of $\mathcal{E}$. Again, similarly to what was observed for KdV in the previous example, by means of $v^{2}=v_{x}^{3}/\lambda$ and $v^{1}=\lambda e^{-z}v_{t}^{3}$ the first-order linear system can be reduced to a higher-order linear one
$\left\{\begin{array}[]{l}\phi_{xx}=z_{x}\phi_{x}+\lambda^{3}e^{-z}\phi_{t},\vskip 5.0pt\\ \phi_{xt}=e^{z}\phi,\vskip 5.0pt\\ {\displaystyle\phi_{tt}=z_{t}\phi_{t}+\frac{1}{\lambda^{3}}e^{-z}\phi_{x}},\end{array}\right$. (29)
where $\phi:=v^{3}$. Indeed, one can readily see that $\mathcal{E}$ is also the integrability condition of (29). Moreover, similarly to the previous example, the $3$-dimensional covering defined by (28) is the extension of a $2$-dimensional covering, by means of the equations for the potential of a conservation law. Indeed, by introducing $\rho^{1}=v^{1}/v^{3}$ and $\rho^{2}=v^{2}/v^{3}$ one can readily see that (28) reduces to
$\left\{\begin{array}[]{l}\left(\ln v^{3}\right)_{x}=\lambda{\small\rho^{2}},\vskip 5.0pt\\ {\displaystyle\left(\ln v^{3}\right)_{t}=\frac{1}{\lambda}e^{z}{\small\rho^{1}}},\vskip 5.0pt\\ \rho_{x}^{1}=-\lambda\,\rho^{1}\rho^{2}-z_{x}\,\rho^{1}+\lambda,\vskip 5.0pt\\ {\displaystyle\rho_{t}^{1}=-\frac{1}{\lambda}{{\rm e}^{z}}{\left(\rho^{1}\right)}^{2}+\frac{1}{\lambda}e^{-2z}\rho^{2}},\vskip 5.0pt\\ \rho_{x}^{2}=-\lambda{\left(\rho^{2}\right)}^{2}+z_{x}{\rho^{2}}+\lambda\rho^{1},\vskip 5.0pt\\ {\displaystyle\rho_{t}^{2}=-\frac{1}{\lambda}{{\rm e}^{z}}{\rho^{1}}{\rho^{2}}+\frac{1}{\lambda}e^{z}},\vskip 5.0pt\\ z_{xt}=-{{\rm e}^{-2\,z}}+{{\rm e}^{z}}.\end{array}\right$.
Thus (28) describes an extension of the $2$-dimensional differentiable covering $\tau:\mathcal{Y}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$, defined by the system
$\mathcal{Y}:\quad\left\{\begin{array}[]{l}\rho_{x}^{1}=-\lambda\,\rho^{1}\rho^{2}-z_{x}\,\rho^{1}+\lambda,\vskip 5.0pt\\ {\displaystyle\rho_{t}^{1}=-\frac{1}{\lambda}{{\rm e}^{z}}{\left(\rho^{1}\right)}^{2}+\frac{1}{\lambda}e^{-2z}\rho^{2}},\vskip 5.0pt\\ \rho_{x}^{2}=-\lambda{\left(\rho^{2}\right)}^{2}+z_{x}{\rho^{2}}+\lambda\rho^{1},\vskip 5.0pt\\ {\displaystyle\rho_{t}^{2}=-\frac{1}{\lambda}{{\rm e}^{z}}{\rho^{1}}{\rho^{2}}+\frac{1}{\lambda}e^{z}},\vskip 5.0pt\\ z_{xt}=-{{\rm e}^{-2\,z}}+{{\rm e}^{z}},\end{array}\right$. (30)
by means of the two differential equations $\left(\ln v^{3}\right)_{x}=\lambda{\small\rho^{2}}$ and $\left(\ln v^{3}\right)_{t}=e^{z}{\small\rho^{1}/\lambda}$, for the potential $\ln v^{3}$ of the conservation law $\mu:=\lambda{\small\rho^{2}}\,dx+\frac{1}{\lambda}e^{z}{\small\rho^{1}}\,dt$, defined in $\tau$.
@@
Let $\mathcal{E}\subset J^{k}(\pi)$ and $\mathcal{E}^{\prime}\subset J^{k^{\prime}}(\pi^{\prime})$ be two formally integrable equations, with $\pi:E\rightarrow M$ and $\pi^{\prime}:E^{\prime}\rightarrow M^{\prime}$ such that $\dim M=\dim M^{\prime}$. A Bäcklund transformation from $\mathcal{E}$ to $\mathcal{E}^{\prime}$ is a regular $\mathcal{C}$-morphism $\mathcal{B}:J^{\infty}(\pi)\rightarrow J^{\infty}(\pi^{\prime})$ from $\mathcal{V}$ to $\mathcal{E}^{\prime}$, for some differentiable covering $\tau:\mathcal{V}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$, with $\mathcal{V}\subset J^{l}(\pi)$. In particular $\mathcal{B}$ is an auto-Bäcklund transformation whenever $\mathcal{E}^{\prime}$ is a copy of $\mathcal{E}$.
According to this definition, a Bäcklund transformation can be seen as a pair of $\mathcal{C}$-morphisms
[…]
which is analogous to the point of view followed by Krasil’shchik and Vinogradov in [26]. However, in this article, we will not limit ourselves to considering Bäcklund transformations that keep the independent variables unchanged.
@@
Continuing Example 16 we can provide an example of auto-Bäcklund transformation for KdV. Indeed, by considering the trivial bundle $\pi:\mathbb{R}^{2}\times\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$, $(x,t,z,\rho)\mapsto(x,t)$ and the $1$-dimensional covering $\tau:\mathcal{Y}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$, it can be checked that the $\mathcal{C}$-morphism $\mathcal{B}:J^{\infty}(\pi)\rightarrow J^{\infty}(\pi^{\prime})$ defined by
@@
through the prolongation formulas (5), defines a $\mathcal{C}$-morphism from $\mathcal{Y}$ to $\mathcal{E}^{\prime}=\{z^{\prime}_{t}+6z^{\prime}z^{\prime}_{x}+z^{\prime}_{xxx}=0\}$, which is a copy of $\mathcal{E}$. Hence (31) provides an auto–Bäcklund transformation of KdV. In practice, for any known solution $f_{0}$ of the KdV, such a transformation associates a new solution
@@
where $g_{0}$ is the solution of the Riccati-type system corresponding to $f_{0}$ for some $\lambda_{0}$, i.e.,
$\left\{\begin{array}[]{l}g_{0x}=-g_{0}^{2}-f_{0}-\lambda_{0},\vskip 5.0pt\\ g_{0t}=-2(2\lambda_{0}-f_{0})g_{0}^{2}-2f_{0x}g_{0}+f_{0xx}+2f_{0}^{2}-2\lambda_{0}f_{0}-4\lambda_{0}^{2},\vskip 5.0pt\end{array}\right$.
@@
$f_{i}=-f_{i-1}-2g_{i-1}^{2}-2\lambda_{i-1},$
where each $g_{i-1}$ is the solution of the Riccati-type system corresponding to $f_{i-1}$ and $\lambda_{i-1}$.
@@
Since $\rho=\phi_{x}/\phi$, one can readily check that in terms of the Lax potential above auto-Bäcklund transformation can be written as $z^{\prime}=-z-2(\phi_{x}/\phi)^{2}-2\lambda$, or equivalently as $z^{\prime}=z+2\left(\ln\phi\right)_{xx}$, in view of (26). Thus, for any known solution $f_{0}$ of the KdV, the transformation associates the new solution $f_{1}=f_{0}+2\left(\ln g_{0}\right)_{xx}$ where $g_{0}$ is the corresponding solution of Lax system
$\left\{\begin{array}[]{l}g_{0xx}=-f_{0}g_{0}-\lambda_{0}g_{0},\vskip 5.0pt\\ g_{0t}=2(2\lambda_{0}-f_{0})g_{0x}+f_{0x}g_{0}.\vskip 5.0pt\end{array}\right$.
@@
If in the previous example we augment the space by passing to the bundle $\hat{\pi}:\mathbb{R}^{2}\times\mathbb{R}^{3}\rightarrow\mathbb{R}^{2},(x,t,z,\rho,\hat{\rho})\mapsto(x,t)$, and instead of (27) consider the system
$\hat{\mathcal{Y}}:\quad\left\{\begin{array}[]{l}\hat{\rho}_{x}=-\hat{\rho}^{2}-z-\hat{\lambda},\vskip 5.0pt\\ \hat{\rho}_{t}=-2(2\hat{\lambda}-z)\hat{\rho}^{2}-2z_{x}\hat{\rho}+z_{xx}+2z^{2}-2\hat{\lambda}z-4\hat{\lambda}^{2},\vskip 5.0pt\\ \rho_{x}=-\rho^{2}-z-\lambda,\vskip 5.0pt\\ \rho_{t}=-2(2\lambda-z)\rho^{2}-2z_{x}\rho+z_{xx}+2z^{2}-2\lambda z-4\lambda^{2},\vskip 5.0pt\\ z_{t}=-6zz_{x}-z_{xxx},\end{array}\right$. (32)
obtained by doubling the Riccati-type system, i.e., by augmenting another copy of the Riccati-type system with spectral parameter $\hat{\lambda}\neq\lambda$, then the natural projection $\hat{\tau}:\hat{\mathcal{Y}}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$ is another covering of KdV and the $\mathcal{C}$-morphism $J^{\infty}(\hat{\pi})\rightarrow J^{\infty}(\pi^{\prime})$ defined by
@@
is a $\mathcal{C}$-morphism from $\hat{\mathcal{Y}}$ to the system
$\mathcal{E}^{\prime}:\quad\left\{\begin{array}[]{l}\rho^{\prime}_{x}=-\rho^{\prime}{}^{2}-z^{\prime}-\lambda,\vskip 5.0pt\\ \rho^{\prime}_{t}=-2(2\lambda-z^{\prime})\rho^{\prime}{}^{2}-2z^{\prime}_{x}\rho^{\prime}+z^{\prime}_{xx}+2z^{\prime}{}^{2}-2\lambda z^{\prime}-4\lambda^{2},\vskip 5.0pt\\ z^{\prime}_{t}=-6z^{\prime}z^{\prime}_{x}-z^{\prime}_{xxx}.\end{array}\right$.
In this case, to obtain a new solution $z^{\prime}$ of KdV out of a given one $z=f_{0}$, the first step is to compute the corresponding solution $\rho=g_{0}$ of the Riccati-type system, with arbitrary $\lambda$. Then, since $g_{0}$ depends on $\lambda$, one gets another instance $\hat{g}_{0}$ of that function, by substituting $\lambda$ with $\hat{\lambda}\neq\lambda$. This way (33) provides the new pair of functions
$z^{\prime}=-f_{0}-2\hat{g}_{0}^{2}-2\hat{\lambda},\qquad\rho^{\prime}=-\frac{\hat{\lambda}-\lambda}{\hat{g}_{0}-g}-\hat{g}_{0},$
@@
$\left\{\begin{array}[]{l}\rho_{x}=-\rho^{2}-z-\lambda,\vskip 5.0pt\\ \rho_{t}=-2(2\lambda-z)\rho^{2}-2z_{x}\rho+z_{xx}+2z^{2}-2\lambda z-4\lambda^{2},\vskip 5.0pt\\ z_{t}=-6zz_{x}-z_{xxx}.\end{array}\right$.
The peculiarity of this transformation, in comparison with that of previous example, is that it provides the new solution $z^{\prime}$ together with the corresponding solution $\rho^{\prime}$ of the Riccati-type system, with arbitrary $\lambda$. Thus, further applications of the transformations only require derivations and algebraic operations.
@@
which is known in literature as the Darboux transformation of KdV (see for instance [22, 33]). Thus, the notion of Bäcklund transformation described by Definition 18 allows one to regard also Darboux transformations as examples of Bäcklund transformations.
@@
Consider the covering $\tau:\mathcal{Y}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$ of the short-pulse equation $\mathcal{E}=\{z_{xt}=\frac{1}{2}z^{2}z_{xx}+zz_{x}^{2}+z\}\subset J^{3}(\pi)$, defined by the system
$\mathcal{Y}:\quad\left\{\begin{array}[]{l}{\displaystyle\rho_{x}=-{\displaystyle\frac{z_{x}}{4\eta}}\rho^{2}+{\displaystyle\frac{1}{2\eta}}\,\rho+{\displaystyle\frac{z_{x}}{4\eta}}},\vskip 10.0pt\\ {\displaystyle\rho_{t}=-{\displaystyle\frac{z\left(zz_{x}+4\,\eta\right)}{8\eta}}\rho^{2}+{\displaystyle\frac{\left(8\,\eta^{2}+z^{2}\right)}{4\eta}}\rho-{\displaystyle\frac{z\left(-zz_{x}+4\,\eta\right)}{8\eta}}},\vskip 10.0pt\\ {\displaystyle z_{xt}={\displaystyle\frac{1}{2}}z^{2}z_{xx}+zz_{x}^{2}+z},\end{array}\right$. (34)
where $\pi:\mathbb{R}^{2}\times\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$, $(x,t,z,\rho)\mapsto(x,t)$. One can check that the transformation
$x^{\prime}=x+8\frac{\eta}{\rho^{2}+1},\qquad t^{\prime}=t,\qquad z^{\prime}=-z+8\eta\frac{\rho}{\rho^{2}+1},$ (35)
determines an auto-Bäcklund transformation, i.e., a Bäcklund transformation from $\mathcal{E}=\{z_{xt}=\frac{1}{2}z^{2}z_{xx}+zz_{x}^{2}+z\}$ to $\mathcal{E}^{\prime}=\{z^{\prime}_{xt}=\frac{1}{2}z^{\prime}{}^{2}z^{\prime}_{xx}+z^{\prime}z^{\prime}{}_{x}^{2}+z^{\prime}\}$.